OpenStudy (anonymous):

A doctor prescribes an ointment that is 2% hydrocortisone. A pharmacist has 1% and 5% concentrations in stock. How many ounces of each should the pharmacist use to make a 9-ounce tube?

OpenStudy (anonymous):

you want 9 ounches of 2%. so the total amount of 1% and 5% lotion has to be total to 9. if you add 2ounches of the 1% to 2ounches of the 5% you would get 4 ounches of 3% (the mean value of the two since you add it in equal ammounts. now if you would only do 1 ounch of the 5% and 2 of the 1% you would do: (2* 1 + 1 * 5)/3 =2.3333 percent and a 3 ounch lotion. try to find the values that give you 2% then convert it to 9 ounches

OpenStudy (anonymous):

Let x ounces of 1% and (9-x) of 5% of hydrocortisone be be mixed to form 9 ounces of 2% hydrocortisone. \[x*\frac{ 1 }{100 }+(9-x)*\frac{ 5 }{100 }=9*\frac{ 2 }{100 }\] orx+5(9-x)=18 solve for x ,you get the solution.

OpenStudy (anonymous):

so I come up with 6.75 oz is that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but which concentration does it go to the 5% or the one%

OpenStudy (anonymous):

it goes to 1% as i have mentioned x ounces of 1%

OpenStudy (anonymous):

thanks that helps out a lot!

OpenStudy (anonymous):

yw